Jamie just paid off a loan he took out six months ago at 12% simple annual interest. He paid $3,816.00, which was the sum of the principal and the simple interest accrued over the length of the loan. What was the amount of principal he borrowed?
Give answer in dollars rounded to the nearest cent. Do NOT enter "$" sign in answer.

Answer by stanbon(75887)

Jamie just paid off a loan he took out six months ago at 12% simple annual interest. He paid $3,816.00, which was the sum of the principal and the simple interest accrued over the length of the loan. What was the amount of principal he borrowed?
Give answer in dollars rounded to the nearest cent. Do NOT enter "$" sign in answer.
-----
A(t) = P(1+rt)
3816 = P(1 + 0.12(1/2))
-----
3816 = P(1.06)
---
P = 3816/1.06 = 3600
-------------
Cheers,
Stan H.
--------------